# Meta-reinforcement learning with minimum attention

### v2

@@ Abstract
Keywords: Reinforcement learning, model-based RL, meta-learning, minimum attention
@@ 1 Introduction
Minimum attention considers the changes of control in state and time, first proposed by Brockett (brockett1997), with the following criterion:
@@
highly relevant in emulating biological control, such as motor learning. As a first step for a rigorous analysis of minimum attention, the existence of the minimum attention was shown, focused on time-variant linear feedback control in the context of motor learning (lee2022). Interestingly, a similar framework was considered in the regularization of fluid flow control by Gunzburger. In the flow control of Navier-Stokes equations, the well-posedness of boundary flow control with the changes of control in time and space was proved by the one-shot method (gunzburger2000). The minimum attention implies “regularization” and might work as stabilization in mathematical point of view, and provides a generative formalism of transition between “closed” and “open” loop controls (lee2022) from the control point of view.
In this paper, we apply minimum attention to the reward of reinforcement learning and investigate its connection to meta-learning. The primary motivation is that minimum attention is a generative formalism of transition between feedforward and feedback controls while getting close to the target goals gradually in state and time (lee2022). This is conjectured to be associated with the adaptation to a new environment in the sense of meta-learning. Overall, the specific nature of the contributions is empirical in the following:
@@
This is indeed significant in its connection to safe RL and few shots of meta-learning. Most learning processes are through experiences, and their functionality is tested with new tasks. Deep reinforcement learning (RL) has been successful in showing competence in learning controls for complex systems (mnih2015). Many of the algorithms are model-free and take the learning process purely by observing state transitions without explicit knowledge of internal system dynamics. As a consequence, most of those algorithms require a large number of sampling and immense learning time to come to a converging stage. Model-based RL shows a promising efficiency in sampling in comparison to model-free RL, but in many cases, the converging return is lower than the asymptotic performance of model-free RL due to model bias. Consequently, the main challenges are how to tolerate the inherent model-bias, and how the learned models and control are flexible to diverse tasks in the presence of uncertainties. Here, in our scope of consideration, two major uncertainties are attributed to 1) the model is not a priori known, and evolves in the model learning process, i.e., model-bias and 2) the model has intrinsic stochasticity or in the interaction with the environment. Model-ensemble approach with a number of models learned together shows a potential to overcome part of the issues (clavera2018; kurutach2018). In the paradigm of “learning to learn”, meta-learning pursues adaptation of learning with few shots confronting new tasks (finn2017), showing a promising learning performance coupled to a model-ensemble approach (clavera2018). In our strategy to handle the uncertainty of model learning, we also incorporate an ensemble of model systems in the course of model learning (kurutach2018), utilizing data collected from the real environment.
Figure 1: Meta-learning from model and environmental perturbations. (a) Half-Cheetah with varying body masses (KiminLee2020), (b) Adaptation to uphill and downhill, (c) One hind leg crippled (nagabandi2019)
Overall, related works on model-based RL and meta-RL, and the associated regularizations are covered in Section 2. We formulate the ensemble model-based meta-policy learning with minimum attention in Section 3. In the main results of Section 4, the training/testing of high-dimensional dynamical systems is compared with the state-of-the art model-free and model-based RLs and analyzed in the feedback and feedforward terms and adaptation in model or environmental perturbations.
@@ 2 Related Works
Model-based reinforcement learning Once initially unknown or partially known models are reduced in model bias and uncertainty, and are well identified, model-based RL approaches can populate imaginary exploration and show outperforming learning curves with smaller samples for policy training. Taking a meta-learning with gradient-descent of a finite number model learning, i.e., ensemble model-based RL shows competent performances (kurutach2018; clavera2018). In the efforts of meta-learning and generalization, KiminLee2020 takes the dynamics model as a global one, which is learned by characterizing a context latent vector for the local dynamics and then by taking conditional prediction. du2020 used neural-ODE in a model-based RL for semi-Markov decision processes in continuous-time dynamics.
Reinforcement meta-learning Model-agnostic meta-learning is by two-level gradient-based optimization in the sense of predictor-corrector schemes (finn2017). For a rapid adaptation to a new environment, it is possible to take off-line learning (clavera2018) or online learning (nagabandi2019) at the test stage in model-based reinforcement learning. In the off-policy actor-critic methods, meta-critic is rapidly learned online for a single task, rather than slowly over a family of tasks (WeiZhou2020). In the technique of generalization, a loss function can be extended to encourage the context latent vector to be trained for predicting both forward and backward dynamics (KiminLee2020).
Regularization in reinforcement learning and meta-learning Regularization is important transforming non-convex profiles to convex and sparse discreteness to continuous and differentiable ones. In machine learning, robustness and domain generalization are fundamentally correlated, but ”robustness” is neither necessary nor sufficient for transferability; rather, regularization is a more fundamental perspective for understanding domain transferability (XiaojunXu2022). In the inference of nonlinear state-space models, Corenflos2021 overcome the issue of traditional resampling encountering non-differentiable loss functions in partial filter methods with entropic regularization. Regularization with Jacobian term stabilizes the random and adversarial input perturbations in the classification without severely degrading generalization (hoffman2019). Biased regularization overcomes complexities in distributions of tasks via a conditioning function which does mapping task’s side information into a meta-parameter vector relevant for the specific task at hand (denevi2020).
@@ 3.1 Mathematical formulation
Given the first-order nonlinear system,
$\dot{x}=f(x,u,t),$ (2)
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control, and $f$ is continuously differentiable with respect to $x$ and $u$. The corresponding stochastic differential equation is the following:
@@ 3.2 Model learning
The model $\{\hat{f}_{\theta_{\mathcal{M}}}\}$ parameterized by $\theta_{\mathcal{M}}$ is learned by the following loss function, similar to the approach of the probabilistic dynamics model (clavera2018):
@@ 3.3 Meta-learning with model-ensemble
The control $u=u_{\theta_{u}}(x)$ is defined by a feed-forward neural network with the parameter $\theta_{u}$ and deterministically by the state $x$. The augmented hyperparameter $\theta$ is the following:
$\theta=\left[\begin{array}[]{c}\theta_{u}\\ \theta_{\mathcal{M}}\end{array}\right].$ (8)
The meta-learning by one-step gradient-ascent in policy optimization proceeds by two steps as follows:
@@
where the sample trajectories $\mathcal{T}_{i}$ are generated by model $\mathcal{M}_{i}$} and the control $u_{\theta}$. Based on the vanilla policy gradient method (peters2006):
$\nabla_{\theta_{u}}\mathcal{J}_{i}(\theta)=-\mathbb{E}_{\mathcal{T}_{i}}\left(\sum_{k=0}^{T}\nabla_{\theta_{u}}\log P_{i}(x_{k},t_{k};\theta)(\sum_{l=k}^{T}\gamma^{l}r_{\rm reg}(x_{l},u_{\theta}(x_{l}))-b_{k})\right),$ (11)
@@
$r_{\rm reg}(x,u_{\theta^{\prime}_{i}}(x))$ $=$ $r(x,u_{\theta^{\prime}_{i}}(x))-\alpha(\|\frac{\partial u_{\theta^{\prime}_{i}}}{\partial x}\|^{2}+\|\frac{\partial u_{\theta^{\prime}_{i}}}{\partial t}\|^{2}),$
$\max_{\theta_{u}}$ $\frac{1}{M}\sum_{i}^{M}\mathcal{J}_{i}(\theta^{\prime})\penalty 10000\ \penalty 10000\ \penalty 10000\ {\rm s.t.}\penalty 10000\ \penalty 10000\ \penalty 10000\$(13)
@@ 3.4 Trust Region Policy Optimization (TRPO)
Based on the trust region policy optimization (TRPO) (schulman2015),
$\max_{\theta}\penalty 10000\ [\nabla_{\theta}\mathcal{J}(\theta_{u})\cdot(\theta_{u}^{\prime}-\theta_{u})]$ (14)
${\rm subject\penalty 10000\ to}\penalty 10000\ \frac{1}{2}\|\theta_{u}^{\prime}-\theta_{u}\|^{2}\leq\delta$
where $\nabla_{\theta_{u}}\mathcal{J}_{\theta}$ follows Vanilla Gradient Policy (VPG) in Equation (11).
3.5 Formulation of control: linearization with stochastic terms
where the stochastic term of control, $\epsilon$ is formulated by Gaussian noise or Ornstein-Uhlenbeck process (uhlenbeck_ornstein):
Algorithm 1 Meta-learning with minimum attention
 Initialize $\theta_{i}=[\theta_{\mathcal{M}}^{i}\penalty 10000\ \theta_{u}]$; $\theta_{\mathcal{M}}^{i}$ for model $\mathcal{M}_{i}$} and $\theta_{u}$ for the control.
 The model ensemble $\mathcal{M}$ = $\{{\mathcal{M}_{i}}\}=\{\hat{f}_{i}\}_{i=1}^{M}$
$R_{i}\leftarrow\emptyset$
for $epoch=1$ to $T_{\rm epoch}$ do
   {Model learning}
  for $i=1$ to $M$ do
   Sample trajectories $\{X_{n}^{i}\}$ from the environment with $\theta_{\mathcal{M}}^{i}\penalty 10000\ \theta^{\prime}_{u,i}]$, and $R_{i}\leftarrow\{X_{n}^{i}\}$:
   Train $\mathcal{M}_{i}$} using $R_{i}$ and $\mathcal{L}(\theta_{\mathcal{M}})$ in Eq. (5).
  end for
   {Meta-policy learning}
  for $i=1$ to $M$ do
   Generate $\{x_{j}^{k}\}$ with $x_{j}^{0}\sim\rho_{i}(\cdot,0)$ and the sample size NN.
   This is to sample model-based ”imaginary” trajectories $\mathcal{T}_{i}$ from $\mathcal{M}_{i}$} using $\theta_{i}$.
   $\mathcal{J}_{i}(\theta)\leftarrow\frac{1}{N}\sum_{k,j}r(x_{j}^{k},u_{\theta}(x_{j}^{k},t_{k}))$
   $\mathcal{J}_{i}(\theta)\leftarrow-\frac{\alpha}{N}\sum_{k,j}(\|\frac{\partial u_{\theta}}{\partial x}\|^{2}+\|\frac{\partial u_{\theta}}{\partial t}\|^{2})(x_{j}^{k},t_{k})$
   Adapt parameters $\theta^{\prime}_{u,i}=\theta_{u}+\beta\nabla_{\theta_{u}}\mathcal{J}_{i}(\theta)$ with VPG/SAC.
   Sample model-based trajectories $\mathcal{T}^{\prime}_{i}$ from $\mathcal{M}_{i}$} using adapted $u_{\theta^{\prime}_{i}}$:
   Generate $\{x_{j}^{k}\}$ with $x_{j}^{0}\sim\rho^{\prime}_{i}(\cdot,0)$ and the sample size NN.
   $\mathcal{J}_{i}(\theta^{\prime})\leftarrow\frac{1}{N}\sum_{k,j}r(x_{j}^{k},u_{\theta^{\prime}_{i}}(x_{j}^{k},t_{k}))$
   $\mathcal{J}_{i}(\theta^{\prime})\leftarrow-\frac{\alpha}{N}\sum_{k,j}(\|\frac{\partial u_{\theta^{\prime}}}{\partial x}\|^{2}+\|\frac{\partial u_{\theta^{\prime}}}{\partial t}\|^{2})(x_{j}^{k},t_{k})$
  end for
  Train $\theta_{u}$ using $\frac{1}{M}\sum_{i}\mathcal{J}_{i}(\theta^{\prime})$ in Eq. (13) with TRPO.
end for
@@ 4 Main Results and Discussion
We conduct a series of experiments across several MuJoCo environments to rationalize the proposed method. The evaluation is designed to demonstrate that the minimum attention regularization: 1) improves asymptotic performance and learning stability; 2) induces policies that are smoother and more energy-efficient by learning a more structured control strategy; and 3) enhances an agent’s ability to rapidly adapt to out-of-distribution (OOD) perturbations.
We present results for HalfCheetah in detail, followed by summaries for Hopper, Walker2d, and Humanoid to demonstrate the generality of our findings.
@@ 4.1 Experimental setup: high-dimensional dynamical systems
Our method integrates the minimum attention regularization into the actor-critic updates of a state-of-the-art model-based meta-learning framework, MB-MPO (clavera2018), which uses SAC (haarnoja2018) for policy optimization. We compare against the original MB-MPO and a purely model-free SAC as baselines.
@@
We mainly explore motor learning and control of four agents of MuJoCo in the OpenAI Gym (brockman2016openai): Half-Cheetah (17-dim states and 6-dim controls), Walker2D (17-dim states and 6-dim control), Hopper (11-dim states and 3-dim controls), and HumanoidTrucated (45-dim states and 17-dim controls). The governing equations for MuJoCo dynamics are formulated by Eqs. (3, 15, 16). The involved learning processes are alternated between model learning and meta-policy learning, and with or without regularization of minimum attention laws. In regard to hyperparameters, key hyperparameters for the Model-Based Meta-Policy Optimization (MB-MPO) with minimum attention implementation and the baseline are described in Appendix A. In general evaluation, we have used 10 random seeds for training episodes. For the final evaluation of a trained agent, we run 10 episodes per seed, i.e., with the randomized initial configuration. Unless otherwise specified, results are averaged over 10 random seeds, with error bands and $\pm$ values indicating one standard deviation. A comprehensive description of all network architectures and hyperparameters is provided in Appendix A. Average total reward, standard deviation/error of reward across seeds/episodes are shown in Figure 2 and 3.
In the meta-training, to analyze the effect of attention weight, we trained agents with attention weight values $\alpha\in\{0,0.01,0.05,1.0,5.0\}$ while keeping all other hyperparameters same. In order to analyze the effects of the number of ensemble models, $M$, we trained 4 different sizes of $M\in\{1,5,15,20\}$. To characterize the effect of the imaginary trajectory size $N$ in the MB-MPO inner loop, we trained the motor learning with 4 different values of $N\in\{16,128,256,512\}$ (Figure 4).
@@
In regards to specific metrics, 1) average feedback norm: during evaluation, we run the trained agent for 10 evaluation episodes and average $\partial u/\partial x||^{2}$ per step, 2) average feedforward norm: define as the averaged $\partial u/\partial t||^{2}$ over 10 evaluation episodes, 3) average energy: define total sum $u^{2}||$ per step as ``control cost". Their time courses are shown in Figure 5. Heatmaps of the Jacobian norm (Figure 6, Figure S1, S4, S7) and jerk norm (Figure 7, Figure S2, S5, S8) between the baselines are shown to qualitatively judge the impact of minimum attention regularization.
@@
In the meta-testing, experiments focus on out-of-distribution tasks and fast adaptation during meta-loop test time evaluation. For each of these experiments, we note the total reward, standard deviation between runs, average feedback, average feedforward norm, and average used energy. With the meta-trained agent, we have evaluated the impact of minimum attention regularization on the learning process by conducting qualitative analysis by visualizing policy sensitivity across different observation of the state with a specific dimension of the environment. The goal is to pick the dimensions that are critical for their respective locomotion and balance challenges. Specifically, we have picked Half-Cheetah, and three projections are plotted by heatmap: Torso height versus forward velocity, Forward velocity versus vertical velocity, and Forward velocity versus angular velocity. Similarly, the projected dimensions for other agents are specified in Appendix A.

### v3

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Minimum attention considers the changes of control in state and time, first proposed by Brockett (Brockett, 1997), with the following criterion:
@@
highly relevant in emulating biological control, such as motor learning. As a first step for a rigorous analysis of minimum attention, the existence of the minimum attention was shown, focused on time-variant linear feedback control in the context of motor learning (Lee and Park, 2022). Interestingly, a similar framework was considered in the regularization of fluid flow control by Gunzburger. In the flow control of Navier-Stokes equations, the well-posedness of boundary flow control with the changes of control in time and space was proved by the one-shot method (Gunzburger and Manservisi, 2000). The minimum attention implies “regularization” and might work as stabilization in mathematical point of view, and provides a generative formalism of transition between “closed” and “open” loop controls (Lee and Park, 2022) from the control point of view.
In this paper, we apply minimum attention to the reward of reinforcement learning and investigate its connection to meta-learning. The primary motivation is that minimum attention is a generative formalism of transition between feedforward and feedback controls while getting close to the target goals gradually in state and time (Lee and Park, 2022). This is conjectured to be associated with the adaptation to a new environment in the sense of meta-learning. Overall, the specific nature of the contributions is empirical in the following:
@@
This is indeed significant in its connection to safe RL and few shots of meta-learning. Most learning processes are through experiences, and their functionality is tested with new tasks. Deep reinforcement learning (RL) has been successful in showing competence in learning controls for complex systems (Mnih et al., 2015). Many of the algorithms are model-free and take the learning process purely by observing state transitions without explicit knowledge of internal system dynamics. As a consequence, most of those algorithms require a large number of sampling and immense learning time to come to a converging stage. Model-based RL shows a promising efficiency in sampling in comparison to model-free RL, but in many cases, the converging return is lower than the asymptotic performance of model-free RL due to model bias. Consequently, the main challenges are how to tolerate the inherent model-bias, and how the learned models and control are flexible to diverse tasks in the presence of uncertainties. Here, in our scope of consideration, two major uncertainties are attributed to 1) the model is not a priori known, and evolves in the model learning process, i.e., model-bias and 2) the model has intrinsic stochasticity or in the interaction with the environment. Model-ensemble approach with a number of models learned together shows a potential to overcome part of the issues (Clavera et al., 2018; Kurutach et al., 2018). In the paradigm of “learning to learn”, meta-learning pursues adaptation of learning with few shots confronting new tasks (Finn et al., 2017), showing a promising learning performance coupled to a model-ensemble approach (Clavera et al., 2018). In our strategy to handle the uncertainty of model learning, we also incorporate an ensemble of model systems in the course of model learning (Kurutach et al., 2018), utilizing data collected from the real environment.
Figure 1: Meta-learning from model and environmental perturbations. (a) Half-Cheetah with varying body masses (Lee et al., 2020), (b) Adaptation to uphill and downhill, (c) One hind leg crippled (Nagabandi et al., 2019)
@@ 2 Related Works
Model-based reinforcement learning Once initially unknown or partially known models are reduced in model bias and uncertainty, and are well identified, model-based RL approaches can populate imaginary exploration and show outperforming learning curves with smaller samples for policy training. Taking a meta-learning with gradient-descent of a finite number model learning, i.e., ensemble model-based RL shows competent performances (Kurutach et al., 2018; Clavera et al., 2018). In the efforts of meta-learning and generalization, Lee, et al. takes the dynamics model as a global one, which is learned by characterizing a context latent vector for the local dynamics and then by taking conditional prediction (Lee et al., 2020). Du, et al. used neural-ODE in a model-based RL for semi-Markov decision processes in continuous-time dynamics (Du et al., 2020).
Reinforcement meta-learning Model-agnostic meta-learning is by two-level gradient-based optimization in the sense of predictor-corrector schemes (Finn et al., 2017). For a rapid adaptation to a new environment, it is possible to take off-line learning (Clavera et al., 2018) or online learning (Nagabandi et al., 2019) at the test stage in model-based reinforcement learning. In the off-policy actor-critic methods, meta-critic is rapidly learned online for a single task, rather than slowly over a family of tasks (Zhou et al., 2020). In the technique of generalization, a loss function can be extended to encourage the context latent vector to be trained for predicting both forward and backward dynamics (Lee et al., 2020).
Regularization in reinforcement learning and meta-learning Regularization is important transforming non-convex profiles to convex and sparse discreteness to continuous and differentiable ones. In machine learning, robustness and domain generalization are fundamentally correlated, but ”robustness” is neither necessary nor sufficient for transferability; rather, regularization is a more fundamental perspective for understanding domain transferability (Xu et al., 2022). In the inference of nonlinear state-space models, Corenflos, et al. overcome the issue of traditional resampling encountering non-differentiable loss functions in partial filter methods with entropic regularization (Corenflos et al., 2021). Regularization with Jacobian term stabilizes the random and adversarial input perturbations in the classification without severely degrading generalization (Hoffman et al., 2019). Biased regularization overcomes complexities in distributions of tasks via a conditioning function which does mapping task’s side information into a meta-parameter vector relevant for the specific task at hand (Denevi et al., 2020).
@@ 3.1 Mathematical formulation
Given the first-order nonlinear system, $\dot{x}=f(x,u,t),$ where $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control, and $f$ is continuously differentiable with respect to $x$ and $u$. The corresponding stochastic differential equation is the following:
@@ 3.2 Model learning
The model $\{\hat{f}_{\theta_{\mathcal{M}}}\}$ parameterized by $\theta_{\mathcal{M}}$ is learned by the following loss function, similar to the approach of the probabilistic dynamics model (Clavera et al., 2018):
@@ 3.3 Meta-learning with model-ensemble
where the stochastic term of control, $\epsilon$ is formulated by Gaussian noise or Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein, 1930):
The control hyperparameter $\theta_{u}$ is constituted with the hyperparameters for $K(t)$ and $v(t)$. Overall, the augmented hyperparameter $\theta=\left[\theta_{u}~\theta_{\mathcal{M}}\right]^{T}$. The meta-learning by one-step gradient-ascent in policy optimization proceeds by two steps as follows:
@@
where the sample trajectories $\mathcal{T}_{i}$ are generated by model $\mathcal{M}_{i}$} and the control $u_{\theta}$. Based on the vanilla policy gradient method (Peters and Schaal, 2006):
$\nabla_{\theta_{u}}\mathcal{J}_{i}(\theta)$ $=$ $-\mathbb{E}_{\mathcal{T}_{i}}(\sum_{k=0}^{T}\nabla_{\theta_{u}}\log P_{i}(x_{k},t_{k};\theta)$
$(\sum_{l=k}^{T}\gamma^{l}r_{\rm reg}(x_{l},u_{\theta}(x_{l}))-b_{k})),$
@@
$r_{\rm reg}(x,u_{\theta^{\prime}_{i}}(x))$ $=$ $r(x,u_{\theta^{\prime}_{i}}(x))$
$-\alpha(\|\frac{\partial u_{\theta^{\prime}_{i}}}{\partial x}\|^{2}+\|\frac{\partial u_{\theta^{\prime}_{i}}}{\partial t}\|^{2}),$
$\max_{\theta_{u}}$ $\frac{1}{M}\sum_{i}^{M}\mathcal{J}_{i}(\theta^{\prime})~~{\rm s.t.}$ (11)
@@ 3.4 Trust Region Policy Optimization (TRPO)
Based on the trust region policy optimization (TRPO) (Schulman et al., 2015), Eq. (11) is trained by the following:
$\max_{\theta_{u}}~[\nabla_{\theta_{u}}\mathcal{J}(\theta_{u})\cdot(\theta_{u}^{\prime}-\theta_{u})]$ (12)
${\rm subject~to}~\frac{1}{2}\|\theta_{u}^{\prime}-\theta_{u}\|^{2}\leq\delta$
where $\nabla_{\theta_{u}}\mathcal{J}_{\theta}$ follows Vanilla Gradient Policy (VPG) in Equation (3.3).
@@ 4 Main Results and Discussion
We conduct a series of experiments across several MuJoCo environments to rationalize the proposed method. The evaluation is designed to demonstrate that the minimum attention regularization 1) improves asymptotic performance and learning stability, 2) induces policies that are smoother and more energy-efficient by learning a more structured control strategy, and 3) enhances an agent’s ability to rapidly adapt to out-of-distribution (OOD) perturbations.
We present results for HalfCheetah in detail, followed by summaries for Hopper, Walker2d, and Humanoid in Appendix B to demonstrate the generality of our findings.
@@ 4.1 Experimental setup: high-dimensional dynamical systems
Our method integrates the minimum attention regularization into the actor-critic updates of a state-of-the-art model-based meta-learning framework, MB-MPO (Clavera et al., 2018), which uses SAC (Haarnoja et al., 2018) for policy optimization. We compare against the original MB-MPO and a purely model-free SAC as baselines.
@@
We mainly explore motor learning and control of four agents of MuJoCo in the OpenAI Gym (Brockman et al., 2016): Half-Cheetah (17-dim states and 6-dim controls), Walker2D (17-dim states and 6-dim control), Hopper (11-dim states and 3-dim controls), and HumanoidTrucated (45-dim states and 17-dim controls). The governing equations for MuJoCo dynamics are formulated by Eqs. (2, 5, 6). The involved learning processes are alternated between model learning and meta-policy learning, and with or without regularization of minimum attention laws. In regard to hyperparameters, key hyperparameters for the Model-Based Meta-Policy Optimization (MB-MPO) with minimum attention implementation and the baseline are described in Appendix A. In general evaluation, we have used 10 random seeds for training episodes. For the final evaluation of a trained agent, we run 10 episodes per seed, i.e., with the randomized initial configuration. Unless otherwise specified, results are averaged over 10 random seeds, with error bands and $\pm$ values indicating one standard deviation. A comprehensive description of all network architectures and hyperparameters is provided in Appendix A. Average total reward, standard deviation/error of reward across seeds/episodes are shown in Figure 2 and 3.
@@
In regards to specific metrics, 1) average feedback norm: during evaluation, we run the trained agent for 10 evaluation episodes and average $\partial u/\partial x||^{2}$ per step, 2) average feedforward norm: define as the averaged $\partial u/\partial t||^{2}$ over 10 evaluation episodes, 3) average energy: define total sum $u^{2}||$ per step as ``control cost". Their time courses are shown in Figure 4. Heatmaps of the Jacobian norm (Figure 5, S4, S8, S11) and jerk norm (Figure 6, S5, S9, S12) between the baselines are shown to qualitatively judge the impact of minimum attention regularization.
@@
In the meta-testing, experiments focus on out-of-distribution tasks and fast adaptation during meta-loop test time evaluation. For each of these experiments, we note the total reward, standard deviation between runs, average feedback norm, average feedforward norm, and average used energy. With the meta-trained agent, we have evaluated the impact of minimum attention regularization on the learning process by conducting qualitative analysis by visualizing policy sensitivity across different observation of the state with a specific dimension of the environment. The goal is to pick the dimensions that are critical for their respective locomotion and balance challenges. Specifically, we have picked Half-Cheetah, and three projections are plotted by heatmap: Torso height versus forward velocity, Forward velocity versus vertical velocity, and Forward velocity versus angular velocity. Similarly, the projected dimensions for other agents are specified in Appendix A.